Assuming complete dissociation, what is the pH of a 3.77 mg/L Ba(OH), solution? pH =
Answer #1

Consider 1 L of solution.

Molar mass of Ba(OH)2,

MM = 1*MM(Ba) + 2*MM(O) + 2*MM(H)

= 1*137.3 + 2*16.0 + 2*1.008

= 171.316 g/mol

mass(Ba(OH)2)= 3.77 mg

= 0.00377 g

use:

number of mol of Ba(OH)2,

n = mass of Ba(OH)2/molar mass of Ba(OH)2

=(3.77*10^-3 g)/(1.713*10^2 g/mol)

= 2.201*10^-5 mol

volume , V = 1 L

use:

Molarity,

M = number of mol / volume in L

= 2.201*10^-5/1

= 2.201*10^-5 M

So,

[OH-] = 2[Ba(OH)2]

= 2*2.201*10^-5 M

= 4.402*10^-5 M

use:

pOH = -log [OH-]

= -log (4.402*10^-5)

= 4.3563

use:

PH = 14 - pOH

= 14 - 4.3563

= 9.6437

Answer: 9.64